On July 1, 2018, Pearl Industries sold administrative equipment with a book value of $1,000,000 to its subsidiary, Shiek Shoes, for $800,000. At the date of sale, the equipment had a remaining life of five years. It is being straight-line depreciated on Shiek's books. It is now December 31, 2020, the end of the accounting year, and you are preparing the working paper to consolidate the trial balances of Pearl and Shiek. Shiek still owns the equipment.

a. Prepare the required eliminating entries for this intercompany equipment sale for the December 31, 2020, consolidation working paper.

b. It is now December 31, 2021. Prepare the required eliminating entries for this intercompany equipment transaction for the December 31, 2021 consolidation working paper.

c. Now assume that Shiek sells the equipment to an outside party for $400,000 on January 1, 2022. What is the consolidated gain on the sale of equipment? What is the gain reported by Shiek? Prepare the required eliminating entries for the December 31, 2022 consolidation working paper.
Answer #1

ANSWER

As per the given question,

Ques No. Date Account titles Debit Credit
a 31-12-20 Investment in Sheik 140,000
Equipment, net 140,000
To eliminate the beginning-of-year unconfirmed gain.
$200,000 – [($200,000/5) x 1 ½] = $140,000.
Equipment, net 40,000
Depreciation expense 40,000
To eliminate the excess depreciation recorded by Shiek in 2020 ($200,000/5).
b 31-12-21 Investment in Shiek 100,000
Equipment, net 100,000
To eliminate the beginning-of-year unconfirmed gain.
$200,000 – [($200,000/5) x 2 ½] = $100,000.
Equipment, net 40,000
Depreciation expense 40,000
To eliminate the excess depreciation recorded by Shiek in 2020 ($200,000/5).
c 31-12-22 Investment in Shiek 60,000
Gain on sale of equipment 60,000
To recognize the remaining unconfirmed gain as confirmed.   
$200,000 – [($200,000/5) x 3½] = $60,000.

THANK YOU FOR THE QUESTION....KINDLY RATE...IT HELPS ME A LOT